OpenStudy (laviajmani):

Solve the following equation for x 2^7-x * 2^5+4x = 64

OpenStudy (styxer):

You can rewrite 64 as 2^6 and cancel the 2s 7-x +5+4x=6 3x=6-12 x=-6/3 x=-2

OpenStudy (vuriffy):

But doesn't it say x * 2^5?

OpenStudy (vuriffy):

128 - x(32) + 4x = 64. 128 - 28x = 64. -28x = 64. x = -2.29

OpenStudy (styxer):

I did as 2^(7-x) * 2^(5-4x) = 64

OpenStudy (vuriffy):

But you didn't multiple them?

OpenStudy (vuriffy):

2^7-x * 2^5+4x = 64 7 - x * 5 + 4x = 6 7 - 5x + 4x = 6. 7 - x = 6. x = 1.

OpenStudy (vuriffy):

If it was written more clearly, maybe we could have a definite answer.

OpenStudy (styxer):

When you have like 2^2 * 2^3 you sum the powers

OpenStudy (vuriffy):

I never said you were wrong, just not clearly written well enough.

OpenStudy (laviajmani):

Thank you so much @Styxer